OpenStudy (goformit100):

A coin is tossed 1000 times with the following frequencies: Head : 455, Tail : 545 Compute the probability for each event

OpenStudy (goformit100):

@Eyad friend help me

OpenStudy (anonymous):

Lets donate the event of getting :Head by A1 :Tail by A2 P(A1)=455/1000= ? P(A2)=545/1000=?

OpenStudy (goformit100):



OpenStudy (goformit100):

Thank you friend